# Carbon Emission Flow Tracing: Fast Algorithm and California Grid Study

### v1

@@ 3.2 Identifying and Merging Strongly Connected Components
See proof in Appendix 7. Since SCCs do not intersect each other, this guarantees once SCCs are all identified, all buses in the power network are separated into mutually exclusive groups, either as a member of exclusive, non-intersecting SCC or as a standard node. Thus, we define $\tilde{G}^{\text{OPF}}$ as a DAG, where every SCC is grouped and merged as a super bus. In $\tilde{G}^{\text{OPF}}$, each newly-constructed bus $\tilde{i}$ corresponds to the $\text{SCC}_{i}$ of $G^{\text{OPF}}$. For nodes residing outside of SCC in $G^{\text{OPF}}$ and connected to the original SCC, they are now directly connected to the merged new node. For the newly indexed bus $\tilde{i}$, we sum over all generations and loads across the original $\text{SCC}_{i}$ and are denoted as $p_{\tilde{i}}^{g}$ and $p_{\tilde{i}}^{d}$ respectively. The following Proposition guarantees that after the transformation from $G^{\text{OPF}}$ $\tilde{G}^{\text{OPF}}$, the solutions of carbon tracing for nodes outside SCCs are not altered:
@@ 4.1 Network Topology
In order to test the accuracy and efficiency of our carbon tracing algorithm, we utilize the California Test System (CATS) (Taylor et al., 2024), which is a geographically-accurate, synthetic electric grid model without CEII-protected information. The model provides geographic coordinates of the loads and generators, transmission line and transformer parameters, and manually curated grid topology. Fig. 2(a) visualizes the CATS model topology, where transmission line thickness represents voltage level, and the size of the blue dots represents bus load demand. The test power network has 8,870 buses, 2,149 generators, and 10,162 transmission lines. The CATS system is with about 73 GW of existing generation capacity, and 1 GWh of existing energy storage capacity. Compared to IEEE test cases usually with limited number of buses and synthetic load data, due to the nature of CAISO’s large-scale network and CATS realistic loading conditions, benchmarking on this model can provide a comprehensive evaluation of our carbon tracing algorithm.
@@ 4.2 Demand, Generation, and Emission Data
To get the real-time power demand data, we query the CAISO API (CAISO, b) to collect and process the 2019 whole-year data. The original data is in 5 minutes resolution, and we down-sample to hourly demand data for the carbon rate computation. Since CAISO’s published data only include system-level demand, we make the assumption that the load demand distribution remains relatively static. We follow a similar load assignment strategy as the method developed in (Taylor et al., 2024), where the distribution of demand across nodes is determined by historical consumption patterns and demographic factors. Based on CATS’s estimated nodal load vector, we scale the nodal demand at each hour to match the hourly aggregated CAISO system demand data.
@@
In both CATS and CAISO systems, there are no regional or nodal renewable generation values available. To implement our carbon tracing algorithm, we follow the steps described in Appendix 8 to distribute the total CAISO system-level renewable generation to each node based on an optimization-based procedure. The carbon emission rate for renewable generation (i.e., $\gamma_{k}$) is set to be zero in the case study.
@@
Since neither CATS model nor CAISO reveals the import energy price, the power generation imported from other states does not have an estimated cost function which can be applied to solve OPF. To address that, we propose to separate the total import value published by CAISO to each state-bordering node in CATS, and make sure nodal power balance is satisfied after such separation. Once we treat the bordering buses as virtual power generation node with generation equal to import power, we do an estimation for import power’s emission rate. As the total emissions from import can be calculated by subtracting emissions from other generation sources. We then fit the import’s emission rate based on historical $\{{p}^{\text{import}},e^{\text{import}}\}$ data pairs, where ${p}^{\text{import}}$ is the imported power (in MW) and $e^{\text{import}}$ is the imported emission (in metric ton of CO2).
@@
Table 3 in Appendix shows the generator emission rate $\gamma_{k}$ data we used for different types of generators (US Energy Information Administration, 2024). Furthermore, CAISO API (CAISO, a) provides the hourly total emission by each energy source in California at 5 minutes resolution. We treat these aggregated emission values as a reference to validate the locational emission rates and emission per energy resources estimated from our algorithm.
@@ 5.1 Validation of Model Fidelity
Once we derive the GNDF matrix from our carbon tracing algorithm, at each timestep $t$ we can also estimate the system-level total emissions $E$ as
@@
The daily time series comparisons of our estimated $E(t)$ versus CAISO’s reported system-level emissions are shown in Fig. 6, for four randomly selected days in different seasons. We observe that our estimated $E(t)$—based on our proposed nodal power demand estimation and nodal emission rates computation method—closely tracks the CAISO published data on a daily basis. These results indicate that our proposed method can effectively model carbon emission rates first at the nodal level and then aggregate them at the system level.
@@
Table 1: Carbon emission rate (metric tonnes of CO2/MWh) based on the estimated total system-level emission v.s. CAISO published values. Emission rates have units of metric tonnes/MWh.
@@ 5.2 Temporal and Spatial Variations of Emission Rates
It is noteworthy that San Luis Obispo County consistently has the lowest emission rate across all seasons, as the state’s largest single power station—the Diablo Canyon nuclear power plant, with a nameplate capacity of 2,256 MW—is continuously generating carbon-free electricity. Thus, our carbon flow tracing algorithm always analyzes the outflow of low-carbon electricity from this county, marked in dark green in Fig. 7. While the Diablo Canyon power plant is scheduled to retire in a few years, California’s geographical distribution of emission rates may change significantly.
@@
To identify the geographical and temporal disparities in emission profiles, Fig. 8 visualizes the per-capita carbon intensity at 11 a.m. and 11 p.m. on a randomly selected date. The value is calculated by first identifying each county’s total emissions using the locational average emission rate and estimated demand, then dividing by the county’s population. A similar analysis showing LME hourly differences is illustrated in Appendix. We clearly observe that northern Californian counties do not show significant changes in total emissions between daytime and nighttime, mostly likely because their renewable generation comes from hydro and wind rather than solar, exhibiting less daily variations. In contrast, solar contributions in southern counties are more significant during the daytime, leading to a larger difference in per-capita carbon intensity between day and night. In general, there is over 100 times difference in terms of emissions per capita associated with electricity usage, highlighting the geographical disparities. It is also notable that based on the results of aggregated emissions by county, Los Angeles County and San Bernardino County have the highest total emissions due to their large populations and high power demand. Despite that, San Bernardino County exhibits a low per capita carbon emission rate because of a high share of renewable contribution, particularly from solar—thanks to its unique geography and some of the largest solar generation capacities in the state (California Energy Commission, 2024a). Interestingly, major metropolitan areas in California, such as the San Francisco Bay Area, Los Angeles, and San Diego, have high population concentrations. While they do not generate much renewable energy locally, they maintain a moderate per-capita emission profile by importing greener electricity from neighboring counties. Meanwhile, Inyo County and Imperial County have a high ratio of renewable generation to electricity use (California Energy Commission, 2024b).
In Fig. 9, we illustrate how the locational average emission rate vary across all nodes in the 24 hours of June 17th. Each column in the figure represents the occurrences of certain emission rate at a specific bus. There are few outliers with much higher emission rates ($>0.5$ metric tonnes CO2/MWh), which indicate the very infrequent dispatch by coal-fired generators in the CAISO network. Notably, regions with higher renewable energy generation often coincide with lower emissions, indicating the local effects of integrating sustainable practices within the power grids. Coastal regions typically have lower locational average emission rates due to the more diverse energy mix, while inland counties’ emission rates are on average higher.
6 Conclusion and Discussion
In this work, we propose a systematic framework for computing locational carbon emission rates through power flow analysis. Our method supports full ACOPF formulations, enabled by cycle flow identification and elimination techniques. The algorithm achieves high computational efficiency, making it suitable for large-scale, realistic power system applications. We apply this framework to perform a real-world-scale, detailed emission analysis on the California grid. Using the CATS model with realistic system topology and parameters, We develop a practical method for estimating nodal power demand on an hourly basis over the entire year of 2019 to solve the OPF and conduct emission analysis. Our method is firstly validated by comparing the results against CAISO’s published system-level emissions. We then analyze scenarios with and without renewable generation, highlighting the critical role of different generation sources in shaping regional emission patterns. The results reveal notable spatial and temporal disparities in California: inland counties tend to exhibit higher carbon intensity, while coastal and high-renewable regions show lower per-capita emissions. These insights underscore the need for geographically informed policy and infrastructure planning to support effective and equitable decarbonization.
Since we directly imported total system-wise renewable generation and load demand from CAISO in this paper, future work can try to predict them based on demand forecasts and generation prediction. Our paper demonstrates the process of calculating average nodal emission rate when the complete network model is accessible. However, there are instances where the network model might not be available to the interested party. An intriguing research area involves understanding how to estimate the topology, parameters, and emission profiles of the electricity network using publicly available data.

### v2

@@ 3.2 Identifying and Merging Strongly Connected Components
See proof in Appendix A of our online preprint (Shen et al., 2025). Since SCCs do not intersect each other, this guarantees once SCCs are all identified, all buses in the power network are separated into mutually exclusive groups, either as a member of exclusive, non-intersecting SCC or as a standard node. Thus, we define $\tilde{G}^{\text{OPF}}$ as a DAG, where every SCC is grouped and merged as a super bus. In $\tilde{G}^{\text{OPF}}$, each newly-constructed bus $\tilde{i}$ corresponds to the $\text{SCC}_{i}$ of $G^{\text{OPF}}$. For nodes residing outside of SCC in $G^{\text{OPF}}$ and connected to the original SCC, they are now directly connected to the merged new node. For the newly indexed bus $\tilde{i}$, we sum over all generations and loads across the original $\text{SCC}_{i}$ and are denoted as $p_{\tilde{i}}^{g}$ and $p_{\tilde{i}}^{d}$ respectively. The following Proposition guarantees that after the transformation from $G^{\text{OPF}}$ $\tilde{G}^{\text{OPF}}$, the solutions of carbon tracing for nodes outside SCCs are not altered:
@@ 4.1 Network Topology
In order to test the accuracy and efficiency of our carbon tracing algorithm, we utilize the California Test System (CATS) Taylor et al. (2024), which is a geographically-accurate, synthetic electric grid model without CEII-protected information. The model provides geographic coordinates of the loads and generators, transmission line and transformer parameters, and manually curated grid topology. Fig. 2(a) visualizes the CATS model topology, where transmission line thickness represents voltage level, and the size of the blue dots represents bus load demand. The test power network has 8,870 buses, 2,149 generators, and 10,162 transmission lines. The CATS system is with about 73 GW of existing generation capacity, and 1 GWh of existing energy storage capacity. Compared to IEEE test cases usually with limited number of buses and synthetic load data, due to the nature of CAISO’s large-scale network and CATS realistic loading conditions, benchmarking on this model can provide a comprehensive evaluation of our carbon tracing algorithm.
@@ 4.2 Demand, Generation, and Emission Data
To get the real-time power demand data, we query the CAISO API CAISO (b) to collect and process the 2019 whole-year data. The original data is in 5 minutes resolution, and we down-sample to hourly demand data for the carbon rate computation. Since CAISO’s published data only include system-level demand, we make the assumption that the load demand distribution remains relatively static. We follow a similar load assignment strategy as the method developed in Taylor et al. (2024), where the distribution of demand across nodes is determined by historical consumption patterns and demographic factors. Based on CATS’s estimated nodal load vector, we scale the nodal demand at each hour to match the hourly aggregated CAISO system demand data.
@@
In both CATS and CAISO systems, there are no regional or nodal renewable generation values available. To implement our carbon tracing algorithm, we follow the steps described in Appendix B to distribute the total CAISO system-level renewable generation to each node based on an optimization-based procedure. The carbon emission rate for renewable generation (i.e., $\gamma_{k}$) is set to be zero in the case study.
@@
Since neither CATS model nor CAISO reveals the import energy price, the power generation imported from other states does not have an estimated cost function which can be applied to solve OPF. To address that, we propose to allocate the total import value published by CAISO to each state-bordering node in CATS, and make sure nodal power balance is satisfied after such separation. Once we treat the bordering buses as virtual power generation node with generation equal to import power, we do an estimation for import power’s emission rate. As the total emissions from import can be calculated by subtracting emissions from other generation sources. We then fit the import’s emission rate based on historical $\{{p}^{\text{import}},e^{\text{import}}\}$ data pairs, where ${p}^{\text{import}}$ is the imported power (in MW) and $e^{\text{import}}$ is the imported emission (in metric ton of CO2).
@@
Table 3 in Appendix shows the generator emission rate $\gamma_{k}$ data we used for different types of generators US Energy Information Administration (2024). Furthermore, CAISO API CAISO (a) provides the hourly total emission by each energy source in California at 5 minutes resolution. We treat these aggregated emission values as a reference to validate the locational emission rates and emission per energy resources estimated from our algorithm.
@@ 5.1 Validation of Model Fidelity
Once we derive the GNDF matrix from our carbon tracing algorithm, at each timestep $t$ we can also estimate the system-level total emissions $E$ in metric tonnes of CO2 as
@@
The daily timeseries comparisons of our estimated $E(t)$ versus CAISO’s reported system-level emissions are shown in Fig. 6, for four randomly selected days in different seasons. We observe that our estimated E(t)E(t)—based on our proposed nodal power demand estimation and nodal emission rates computation method—closely tracks the CAISO published data on a daily basis. These results indicate that our proposed method can effectively model carbon emission rates first at the nodal level and then aggregate them at the system level.
@@
Table 1: Average carbon emission rate (metric tonnes of CO2/MWh) based on the estimated total system-level emission v.s. CAISO published values.
@@ 5.2 Temporal and Spatial Variations of Emission Rates
It is noteworthy that San Luis Obispo County consistently has the lowest emission rate across all seasons, as the state’s largest single power station—the Diablo Canyon nuclear power plant, with a nameplate capacity of 2,256 MW—is continuously generating carbon-free electricity. Thus, our carbon flow tracing algorithm always analyzes the outflow of low-carbon electricity from this county, marked in dark green in Fig. 7. Since the Diablo Canyon power plant is scheduled to retire in a few years, California’s geographical distribution of emission rates may change significantly. Fig. 8 illustrates how the locational average emission rate vary across all nodes in the 24 hours of June 17th. Each column in the figure represents the occurrences of certain emission rate at a specific bus. There are few outliers with much higher emission rates (>0.5>0.5 metric tonnes CO2/MWh), indicating the very infrequent dispatch by coal-fired generators in the CAISO network. Notably, regions with higher renewable energy generation often coincide with lower emissions, indicating the local effects of integrating sustainable practices. Coastal regions typically have lower LAE rates due to the more diverse energy mix.
Figure 9: (a) Geographical illustration of county-level LME; (b) Distribution of LME rates for all counties in California. Results are shown for 11am, June 17th, 2019.
In Fig. 9 we illustrate the geographical variations of LME at 11am on June 17th, 2019. The colorbar is coded based on double-sided logscale axis. We can observe the LME rates vary significantly across the state. Plumas County in Northeastern California records the highest LME rate of $10.21$ metric tonnes CO2/MWh. This may be due to the relatively low power consumption with small population of $18,807$ in Plumas County, so that the power dispatch is varying a lot depending on loading conditions. LME can also drop below 0, as increasing the total load may lead to more renewables being dispatched, causing the negative marginal emission rate. Such a spatial heterogeneity also underscores the importance of energy management strategies considering distributed energy resources, and provide localized insights for emission reduction initiatives. More results are in Appendix C.
6 Conclusion and Discussion
In this work, we propose a systematic framework for computing locational carbon emission rates through power flow analysis. Our method supports full ACOPF formulations, enabled by cycle flow identification and elimination techniques. The algorithm achieves high computational efficiency, making it suitable for large-scale, realistic power system applications. We apply this framework to perform an ISO-scale, hourly emission analysis on the California grid. Results underscore the need for geographically informed policy and infrastructure planning to support effective and equitable decarbonization.
Future work can incorporate demand forecasts and generation prediction for emission estimation. There are instances where the network model might not be available to the interested party. An intriguing research area involves understanding how to estimate the topology, parameters, and emission profiles of the electricity network using publicly available data.
Appendix A Proof of Proposition 1
Proposition 2.
Sets of SCCs do not intersect each other.
Proof.
[…]
Since $\mathcal{S}_{1}\in\mathcal{S}_{m}$ and $\mathcal{S}_{2}\in\mathcal{S}_{m}$, $|\mathcal{S}_{m}|>|\mathcal{S}_{1}|$ and $|\mathcal{S}_{m}|>|\mathcal{S}_{2}|$. $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are not maximal set of load buses that are any two buses are reachable, i.e. they are not SCCs by definition. So this is proved by contradiction. ∎
Appendix B Data Processing for
Renewable Generation and Imports
As CAISO only publishes aggregated demand and net imports, while in order to calculate directed power flow and associated emission flows, we need to know each node’s renewable power generation. Another challenge is that in the original value in the Matpower file provided by the CATS model (Taylor et al., 2024), import values is not recorded and they will be 0 if directly loaded into OPF solution process. Moreover, there is no cost vector for renewables, making it hard to directly find the market clearing results. To avoid that, we develop a tailored approach to estimate nodal renewable generation based on the total renewables generation published by CAISO.
Let $\mathbf{p}_{total}^{d}$ represent total load vector across all buses under which all types of generators are considered. Let $\mathbf{p}_{total}^{g}$ represent total power generation of all types of generators, and we have $\sum p_{total}^{d}=\sum p_{total}^{g}$ always hold. Under the case where renewable generation is not taken into account, we use $\mathbf{p}_{net}^{g}$ to represent power generation of non-renewable generators and $\mathbf{p}_{net}^{d}$ represent the net load of the buses after excluding renewable generation. We are interested in finding each node’s renewable contribution $p_{i}^{d}(p_{r}^{g})$, with $r\in\mathcal{R}$ denoting the index for renewable generator. We then have
$p_{i,total}^{d}=\sum_{k}p_{i}^{d}(p_{k,net}^{g})+p_{i}^{d}(p_{r}^{g}).$ (10)
To properly assign $p_{i}^{d}(p_{k,net}^{g})$, we propose to solve the OPF model with only the non-renewable generation components, using CATS’s model’s published cost functions for each generator. This approach allows us to obtain a baseline scenario against which we can compare the impacts of integrating renewable generation into the network. The optimization problem will focus on minimizing operational costs while satisfying load demand and other operational constraints as listed in (1), and we use $OPF(\cdot)$ to denote such a solution process and the mapping function from load to corresponding generation:
$\mathbf{p}_{net}^{g}=OPF(\mathbf{p}_{net}^{d});$ (11)
After finding the directed power flow based on $\mathbf{p}_{net}^{g}$ and then determining each non-renewable generator’s nodal power contribution, we are ready to find each node’s renewable generation level while satisfying power balance constraints:
$p_{i}^{d}(p_{r}^{g})=p^{d}_{i}-\sum_{k}p_{i}^{d}(p_{k,net}^{g});$ (12)
After getting the proportion of power supplied by renewable generators at each bus, we can also calculate the carbon intensity of the two scenarios with or without considering renewables dispatch. The results shown across this paper are for the scenario with renewables dispatch, which takes the renewables’ emission reduction potentials into account.
Appendix C More Results on California Grid
C.1 Evaluation Metrics:
We adopt MAPE and wMAPE for evaluating the algorithm performance. We take the evaluation of power generation as an example here. Within evaluation time horizon, $p_{t}^{cg}$ is the actual power generation data from CAISO and $p_{t}^{og}$ is the power generation result from the OPF model.
$MAPE=\dfrac{1}{T}\sum_{t=1}^{T}|\dfrac{p_{t}^{cg}-p_{t}^{og}}{p_{t}^{cg}}|;wMAPE=\dfrac{\sum_{t=1}^{T}|p_{t}^{cg}-p_{t}^{og}|}{\sum_{t=1}^{T}|p_{t}^{cg}|}.$ (13)
Similarly, we can replace $p$ with $\delta$ for the evaluation of carbon emission rate in (13). In Fig. 5 and Fig. 6, we use these metrics to verify the effectiveness of our proposed approach on generation mix and system-level emissions respectively.
C.2 Additional Results
For each month, we randomly select one day as a test sample, and calculate the MAPE and wMAPE across 24 hours. The final results can be found in Table 2. It shows a higher error rate of emission rate calculate in May and June. This could be because the unstable nature of solar and hydro renewable generation during summer months. While for both power generation and emission calculation, our approach can closely track the ground truth value, indicating the our underlying OPF engine can well represent the power dispatch process for CAISO grid. This validation procedure makes sure proposed emission tracking algorithm can provide high fidelity results for locational emission analysis.
Table 2: Carbon emission (C) and power generation (G) MAPE (%) rate across 12 months. Ground truth values are from CAISO’s published emission and generation data.
  Months Jan Feb Mar Apr May June July Aug Sept Oct Nov Dec
MAPE(G) 3.00 6.39 8.00 7.35 13.19 11.64 8.46 6.78 7.47 4.85 3.62 1.97
wMAPE(G) 3.05 6.55 7.61 7.89 13.27 11.26 8.71 6.43 7.75 3.51 3.16 1.81
MAPE(C) 6.84 3.35 4.29 8.90 12.23 13.37 11.40 5.10 5.18 13.29 7.03 2.56
wMAPE(C) 6.08 2.58 4.09 10.16 12.79 12.09 10.48 5.16 4.52 12.09 6.93 2.71
In Fig. 10, we plot and compare the distribution of locational average emission rate in three counties: Amador, Sacramento, and San Diego. A clear discrepancy can be observed in the emission profiles among these counties, highlighting the impact of local generations/transmission and geographical factors on emissions. Amador County, characterized by its rural landscape and lower population density, exhibits significantly more stochastic LAE rates compared to the more urbanized areas.
Figure 10: Comparison of distribution of Amador, Sacramento, and San Diego County’s regional averal emission rate.
@@ C.2 Additional Results
To identify the geographical and temporal disparities in emission profiles, Fig. 11 visualizes the per-capita carbon intensity at 11 a.m. and 11 p.m. on a randomly selected date. The value is calculated by first identifying each county’s total emissions using the locational average emission rate and estimated demand, then dividing by the county’s population. We clearly observe that northern Californian counties do not show significant changes in total emissions between daytime and nighttime, mostly likely because their renewable generation comes from hydro and wind rather than solar, exhibiting less daily variations. In contrast, solar contributions in southern counties are more significant during the daytime, leading to a larger difference in per-capita carbon intensity between day and night. In general, there is over 100 times difference in terms of emissions per capita associated with electricity usage, highlighting the geographical disparities. It is also notable that based on the results of aggregated emissions by county, Los Angeles County and San Bernardino County have the highest total emissions due to their large populations and high power demand. Despite that, San Bernardino County exhibits a low per capita carbon emission rate because of a high share of renewable contribution, particularly from solar—thanks to its unique geography and some of the largest solar generation capacities in the state (California Energy Commission, 2024a). Interestingly, major metropolitan areas in California, such as the San Francisco Bay Area, Los Angeles, and San Diego, have high population concentrations. While they do not generate much renewable energy locally, they maintain a moderate per-capita emission profile by importing greener electricity from neighboring counties. Meanwhile, Inyo County and Imperial County have a high ratio of renewable generation to electricity use (California Energy Commission, 2024b).
In Fig. 12, we illustrate the geographical distribution of locational average emission rate across California at 11 a.m. and 11 p.m. on June 17th, 2019. The figure clearly shows the distinct patterns between daytime and nighttime. Interestingly, the inland counties are observing higher emission rate during night. Counties with higher power demand such as San Francisco, San Diego, and Los Angeles counties experience mild emission rate, partially because the energy mix at these counties are more diverse.
Figure 12: Geographical illustration of carbon emission rate at 11 am and 11 pm on a random selected date.
Appendix D Generator Emission Data
In Table 3, we report the emission rate used in this study, which is based on the US Energy Information Administration (EIA)’s report on emissions by plant and by region (US Energy Information Administration, 2024). It is noteworthy currently in our simulation we use EIA’s published, unified emission rate for Import. Future work can look into each specific import connection’s emission rates, which can further improve carbon tracing results.
Table 3: Carbon emission rate by generator fuel type. t/MWh stands for metric tonnes per megawatt-hour.
  Fuel Type Emission Rate (t/MWh) Category
Coal 0.82 Fossil
Petroleum Liquids 0.656 Fossil
Natural Gas 0.44 Fossil
Nuclear 0.0 Low-Carbon
Hydro 0.0 Renewable
Biomass 0.23 Renewable
Wind 0.0 Renewable
Solar 0.0 Renewable
Geothermal 0.038 Renewable
Other/Import 0.43 Mix